Suppose the volume of a cube is numerically equal to half the surface area of the cube. Find the surface area of this cube. Thanks for your help.
Answer by TimothyLamb(4379)

v = xxx
a = 6xx
---
v = a/2
xxx = 6xx/2
xxx = 3xx
x = 3
---
v = 3*3*3
v = 27 cu.units
---
a = 6xx
a = 6*3*3
a = 54 sq.units
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---